OpenStudy (anonymous):

How many moles of chlorine molecules correspond to 1.81 x 1028 molecules of Cl2?

OpenStudy (anonymous):

1.81 x 10^28

OpenStudy (anonymous):

1.09 x 10^6

OpenStudy (anonymous):

\[1.81*10^{28} molecules Cl_2 \frac{1 mol Cl_2}{6.022*10^{23}moleculesCl_2}=3.01*10^{4} mol Cl_2\]